OpenStudy (anonymous):

I need the steps this one also please (4x+1)(2x^2+x+1)

OpenStudy (anonymous):

use the distributive rule: a(b+c) = ab + ac in this case (4x+1)(2x^2+x+1) = (4x)(2x^2) + (4x)x + (4x)1 + 1(2x^2 + ... Do you get the idea? can you finish this off?